7. Given the following function TR(x)--3x2 + 260x TC(x) = 5,500 + 50x a. Find the revenue maximizing level of output (check 2a order conditions also). b. Find the maximum revenue c. Find the corresponding profit function d. Find the profit maximizing level of output (check 2nd order conditions also) e. Find the maximum profit
Answer #1

a) Revenue is maximized where d(TR)/dx = 0 and d2(TR)/dx2 should be negative (Second order condition)

d(TR)/dx = -6x + 260 = 0

x = 43.33 (output is maximized at this level)

d2(TR)/dx2 = -6 (second order condition is satisfied)

b) TR (max) = - 3 x (43.33)2 + 260 x 43.33 = 5633.33

c) Profit (P) = TR - TC = -3x2 + 210x - 5500

d) Profit is maximized where d(P)/dx = 0 and d2(P)/dx2 should be negative (Second order condition)

dP/dx = -6x + 210 = 0

x = 35 (profit is maximized at this level)

d2P/dx2 = -6 (second order condition is satisfied)

e) Maximum profit = - 1825